two free protons are separated by a distance of 1m. if they are released , the kinetic energy of each proton when at infinite separation is

Grade:12

2 Answers

KARAN ACHARYA
12 years ago

initial energy = kqq/r =kqq        (as r= 1m)

final energy = kinetic energy + potential energy 

                  = kinetic energy  + 0      (as ptential energy  of system is 0 at infinite distance)

 

ENERGY CONSERVATION

final kinetic energy of system =  initial potentil energy of system 

                                          =kqq

hence kinetic energy of each proton = kqq/2

q is charge of proton and value of k  is 9 X (10^9) 

 

 

PLEASE          APPROVE          MY          ANSWER          IF          YOU          LIKED          IT

Chetan Mandayam Nayakar
12 years ago

at 1m separation potential energy =E= (1/4pi(epsilon-zero))e2, at infinite separation potential energy is zero and the two protons have same kinetic energy, equal to E/2,